OpenStudy (milliedelongg):

Need Help ASAP x-y+2z=-2 4x+2y+7z=-9 x+5y+z=-3

jimthompson5910 (jim_thompson5910):

There are a few ways to do this. One way is to use substitution. If you use substitution, then you can solve the first equation `x-y+2z=-2` for x What do you get when you do this?

OpenStudy (milliedelongg):

how am i able to substitute if none of the varibes are the same?

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean by none of the variables being the same?

jimthompson5910 (jim_thompson5910):

are you able to solve x-y+2z=-2 for x?

OpenStudy (milliedelongg):

would it be x=-2+y-2z?

jimthompson5910 (jim_thompson5910):

yes that's correct

jimthompson5910 (jim_thompson5910):

now let's move onto the second equation 4x+2y+7z=-9

jimthompson5910 (jim_thompson5910):

We will replace the 'x' with -2+y-2z to go from 4x+2y+7z=-9 to 4(-2+y-2z)+2y+7z=-9 does this make sense?

OpenStudy (milliedelongg):

yes it does!

OpenStudy (milliedelongg):

now I factor?

jimthompson5910 (jim_thompson5910):

so let's do a bit of simplification after replacing x with -2+y-2z

jimthompson5910 (jim_thompson5910):

4(-2+y-2z)+2y+7z=-9 4(-2)+4(y)+4(-2z)+2y+7z=-9 -8+4y-8z+2y+7z = -9 -8+4y+2y-8z+7z = -9 -8+6y-z = -9 -8+6y-z+8 = -9+8 6y - z = -1 Do you agree with those steps shown above?

OpenStudy (milliedelongg):

I don't understand how you got -8+6y-=-9

jimthompson5910 (jim_thompson5910):

4y+2y = 6y -8z+7z = -1z = -z

jimthompson5910 (jim_thompson5910):

I distributed then combined like terms

OpenStudy (milliedelongg):

wouldn't you subtract though? to set to zero?

jimthompson5910 (jim_thompson5910):

I'm not sure what you mean but all I'm doing in those steps is simplifying the left side. When I get to the second to last step, that's when I add 8 to both sides to move the -8 over

jimthompson5910 (jim_thompson5910):

only focus on `4(-2+y-2z)+2y+7z` I turned `4(-2+y-2z)+2y+7z` into `-8+6y-z` through distributing and combining like terms

OpenStudy (milliedelongg):

okay I understand I just am used to subtracting opposed to adding

jimthompson5910 (jim_thompson5910):

so you see how I got to 6y - z = -1 ?

OpenStudy (milliedelongg):

yes

jimthompson5910 (jim_thompson5910):

now let's move to the equation x+5y+z=-3 (the third given equation at the top)

jimthompson5910 (jim_thompson5910):

x+5y+z=-3 -2+y-2z+5y+z=-3 ... replace x with -2+y-2z what happens when you combine like terms on the left side? What do you get?

OpenStudy (milliedelongg):

-2+6y-1z=-3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we can add 2 to both sides to get... -2+6y-1z=-3 -2+6y-1z+2=-3+2 6y - z = -1 agreed?

OpenStudy (milliedelongg):

agreed c:

jimthompson5910 (jim_thompson5910):

So if we replace x with -2+y-2z in equation B, then we end up with 6y - z = -1 if we replace x with -2+y-2z in equation C, then we end up with 6y - z = -1 Essentially this means that this system is `dependent` and there are infinitely many solutions. All of the solutions will fall on the same line

OpenStudy (milliedelongg):

what would y equal?

jimthompson5910 (jim_thompson5910):

how does your teacher want you to enter the solution format?

OpenStudy (milliedelongg):

I need the answer for x which we got first off and the answer for y

jimthompson5910 (jim_thompson5910):

do you agree how there are infinitely many solutions? is there a special way your teacher wants you to enter the answer? Some teachers will accept "infinite solutions" and you can move on. Other teachers will want you to be more specific

jimthompson5910 (jim_thompson5910):

it will depend on what your teacher wants exactly

OpenStudy (milliedelongg):

I honestly have no idea I have to type in what x and y equal the only other option is that there are no solutions. I entered in what we got for x it was incorrect. the answers were in fractions I am beyond confused it said the answer was x=-11/6z-13/6 and y = 1/6z-1/6

jimthompson5910 (jim_thompson5910):

I see. Ok so here's how they got y \[\Large 6y - z = -1\] \[\Large 6y - z+z = -1+z\] \[\Large 6y = z-1\] \[\Large \frac{6y}{6} = \frac{z-1}{6}\] \[\Large y = \frac{z}{6}-\frac{1}{6}\] \[\Large y = \frac{1}{6}z-\frac{1}{6}\] Does this make sense?

OpenStudy (milliedelongg):

honestly no

jimthompson5910 (jim_thompson5910):

which step are you stuck on?

OpenStudy (milliedelongg):

ugh I honestly don't know can we try again with different numbers? ill try and give you my answers as I go?

jimthompson5910 (jim_thompson5910):

what I did was add 'z' to both sides first then I divided both sides by 6 to isolate y. I broke up the fraction to put it in the form your teacher has

jimthompson5910 (jim_thompson5910):

sure we can try another one if you want

OpenStudy (milliedelongg):

okay the equation is this x-y+4z=-4 2x+2y+7x=-9 x+3y+3z=-5

jimthompson5910 (jim_thompson5910):

what do you get when you isolate x in x-y+4z=-4 ?

OpenStudy (milliedelongg):

I got y-4z-4

jimthompson5910 (jim_thompson5910):

so x = y-4z-4

jimthompson5910 (jim_thompson5910):

now we move to equation B which is 2x+2y+7x=-9 we will replace the 'x' with y-4z-4 and simplify things out 2x+2y+7x=-9 2(y-4z-4)+2y+7x=-9 what do you get when you simplify `2(y-4z-4)+2y+7x` ?

OpenStudy (milliedelongg):

yes and I pluged it into the second equation factored and simplified and got -8+4y-z=-9 then added 8 and got 4y-z=-1

jimthompson5910 (jim_thompson5910):

wait a moment

jimthompson5910 (jim_thompson5910):

you wrote equation B as 2x+2y+7x=-9 is it really supposed to be 2x+2y+7z=-9 ???

OpenStudy (milliedelongg):

yes 7z sorry

jimthompson5910 (jim_thompson5910):

ok I agree. I'm getting 4y-z=-1 as well

jimthompson5910 (jim_thompson5910):

ready to move onto equation C?

OpenStudy (milliedelongg):

yes I got the exact same solution 4y+-z=-1

jimthompson5910 (jim_thompson5910):

ok so equation B and equation C lead to 4y-z = -1

OpenStudy (milliedelongg):

agreed

jimthompson5910 (jim_thompson5910):

since we have two identical equations, it leads to `infinite solutions`

jimthompson5910 (jim_thompson5910):

I'm assuming your teacher wants you to solve for y here?

OpenStudy (milliedelongg):

correct I need for both y and x

jimthompson5910 (jim_thompson5910):